• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ewing/Fansler

    Log entry time 13:59:42 on May 25, 2007

    Entry number 203202

    keyword=Cryo problems

    Last night at approximately 11 PM right Q2 tripped on liquid level. The shift leader called to report the problem. We troubleshot over the phone checking temperatures and valve activity. The magnet recovered and we were able to reset the faults and run again. We were unable to determine if we had a liquid level meter problem, or a cryogenics problem. This morning we investigated the problem further, and determined that the 16 valve card on the right Q1 failed open, causing the Q2 trip we experienced Thursday night. After closer examination, we determined that the liquid level meter for right Q2 was also not giving us an accurate indication of liquid level. We replaced the valve card for the 16 valve on right Q1 and replaced the liquid level meter for right Q2. Both problems were resolved. Q2 ramped normally, but the Q1 displayed multiple faults, and would not respond to a reset command. We cycled power on the Saclay rack and the power supply with no joy. We disconnected the communications cable from the control panel and re-installed…no joy. Mark noticed that the Saclay rack was giving us an indication of a cryogenics fault. The liquid level meter was set to trip at 60%, and the current liquid level was 68%????? We disconnected the analog/relay out cable and the contactor cycled normally. This fixed the problem. We reset the Saclay rack, and power supply. The power supply ramped normally and the cryogenics appear to be stabilizing. We will continue to monitor the cryogenics.

    A copy of this log entry has been emailed to: fansler@jlab.org, stevensm@jlab.org